• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 00:04:28 on May 07, 2008

    Entry number 237480

    keyword=Shift summary (swing)

    16:00 Shift starts with Kinematics N (part III)
    16:33 Asked to increase the beam current to 7uA with raster on (0.6x0.8)
    22:32 Beam stopped to move to Al dummy target
    22:38 Target motion completed. Beam back.Start taking data with Al dummy
    target.
    23:49 Back to LH2 target, continue production data.
    

    Runs 4880 - 4900 Production data with LH2 target Runs 4901, 4902 Al dummy target runs Run 4903 Back to production data with LH2 target